OpenStudy (bbb911):

Porportions 6/7 over 5 equals 10/7 over 3

OpenStudy (anonymous):

(6/7)/5 = ? (10/7)/3 = ? Evaluate: (6/7)/5 = 6/(7*5) = 6/35 (10/7)/3 = ?

OpenStudy (agent0smith):

\[\Huge \frac{ \left( \frac{ 6 }{ 7} \right) }{5 }=\frac{ \left( \frac{ 10 }{ 7 } \right)}{ 3 }\] is this your problem? That would simplify to \[\Large \frac{ 6 }{ 35 } = \frac{ 10 }{ 21 }\] which are not equal.